OpenStudy (anonymous):

A company has 22 staff members - the owner, 3 senior executives, 14 junior executives and 4 support staff. A committee of 8 members of the staff must be convened and the members are selected at random. a) what is the probability that there are fewer support staff on the committee than senior executives. b) what is the probability that the owner in on the committee and there is at least one member from each of the 3 staff categories on the committee too?

OpenStudy (anonymous):

First thing we need to know is how many committees of 8 people can be formed from randomly selecting 8 from 22.

OpenStudy (anonymous):

22C8

OpenStudy (anonymous):

agrreed

OpenStudy (anonymous):

and that would be the n(s) for P(A) =n(A)/n(s) ; A = the event specified in question

OpenStudy (anonymous):

Now, if there are fewer support staff than senior executives this can occur in a number of different ways. 0 support and 1,2 or 3 senior. 1 support and 2 or 3 senior. Or 2 support and 3 senior. This gives n(A)=6.

OpenStudy (anonymous):

So, if my reasoning is correct, you have P(A)=6/22C8

OpenStudy (anonymous):

no...the answer is 81549/319770

OpenStudy (anonymous):

hmmm.

OpenStudy (anonymous):

22C8 = 319770 so i dont know where 81549 comes from

OpenStudy (anonymous):

oh...oooops yeah

OpenStudy (anonymous):

is the question trying to say that with 3 execs, there is the possibility of no support staff or there will be 1 or 2?

OpenStudy (anonymous):

so i think in a case of where there is 3 execs, then no. of ways is 3C3.15C5 + 3C3.4C1.15C4+3C3.4C2.15C3?

OpenStudy (anonymous):

with 3 seniors AND zero support. We should have 15C5. With 3 seniors and 1 support: 16C4. With 3 seniors and 2 support, 17C3. So, 15C5 + 16C4 + 17C3=3003+1820+680=5503. We still have other possibilities to add to this. We could have 2 seniors with zero support. 2 seniors with 1 support. 1 senior with zero support. add them up...

OpenStudy (anonymous):

ah yes...i get the logic now. How would you do the second part of the question?

OpenStudy (anonymous):

hmmmm I still thing I missing something though after reading your last post.

OpenStudy (anonymous):

with 3 seniors and zero support 3C3*15C5=15C5 is correct.

OpenStudy (anonymous):

with three seniors and 1 support we would have 3C3*4C1*16C4

OpenStudy (anonymous):

I got the answer! :D

OpenStudy (anonymous):

sweet...am I on the right track at least? Im rusty on perms/combs/probability

OpenStudy (anonymous):

yip :) thanks to your logic, thats how i got it

OpenStudy (anonymous):

not sure how to do the second part...what would your logic be?

OpenStudy (anonymous):

because the owner must the in the committee, we can only choose from 21 staff

OpenStudy (anonymous):

the number of committees we can form with the owner on it is 21C7

OpenStudy (anonymous):

and from the 7 staff chosen, we allocate their positions in the committee?

OpenStudy (anonymous):

from the seven committee spots left over we need at LEAST one person from each category of employee. there are lots of possibilities ehre

OpenStudy (anonymous):

so like 7C1.6C1.5.1 and then multiply 5C4?

OpenStudy (anonymous):

for the case where we have one from each category

OpenStudy (anonymous):

the amount of 68438 is what we are looking for in n(A)

OpenStudy (anonymous):

well, we pretty much have to itemize each way of doing this

OpenStudy (anonymous):

your teacher is cruel btw

OpenStudy (anonymous):

i very much agree! and he sometimes never completes hard examples with us!

OpenStudy (anonymous):

if we have to itemise everything...this will take so long and i have one more question to ask you (for the day)...should we just give up on that?

OpenStudy (anonymous):

so, one way is 1 owner, 1 senior, 5 junior and 1 support. I get, 1C1*3C1*14C5*4C1

OpenStudy (anonymous):

There must be a shortcut to this one...way too many possibilities

OpenStudy (anonymous):

1 - the probability of the opposite?

OpenStudy (anonymous):

yeah, that's what I was thinking...but...not seeing it

OpenStudy (anonymous):

ok...we need at least 1 member from each group. So the opposite would be having zero members from each group

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

n(A) = 21C7?

OpenStudy (anonymous):

wait, if none of them is chosen...only the owner would be in the group

OpenStudy (anonymous):

so..starting with the seniors group. The probability of forming a committee with zero seniors 19C5. zero seniors and 1 owner: 18C4. At LEAST one senior 1-19C5.

OpenStudy (anonymous):

do u mean 1 - 19C5 - other combinations too?

OpenStudy (anonymous):

Now, the juniors. zero juniors and 1 owner, 7C8=not possible. We MUST have at least one junior on the committee. Thus, probability of having at least one junior =1

OpenStudy (anonymous):

There's an error on my reasoning about the seniors group. The probability of having one owner and at least one senior is 1-P(one owner and zero seniors)=1-(18C7/22C8)

OpenStudy (anonymous):

but we can't have a probability of 1...

OpenStudy (anonymous):

So...the probability of having 1 owner and at least one junior and at least one senior is (1-(18C7/22C8)*1

OpenStudy (anonymous):

now...last group. The prob of having 1 owner and at least one support is 1-P(one owner and zero support)=1-(17C7/22C8)

OpenStudy (anonymous):

then you multiple them all together. overall probability is (1-(18C7/22C8))*(1-(17C7/22C8))

OpenStudy (anonymous):

I'll let you compute this and see if it is correct (if not I think I know where my error likely will be)

OpenStudy (anonymous):

and is not correct. the answer is suppose to be 0.2140

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ur answer gives 0.8457...

OpenStudy (anonymous):

damn....tough one. I think your stuck itemizing them all and adding up the individual possibilities

OpenStudy (anonymous):

wish I could be more help but I have to go to work

OpenStudy (anonymous):

good luck :)

OpenStudy (anonymous):

ok thanks. will u be online tomorrow?

OpenStudy (anonymous):

yeah probably

OpenStudy (anonymous):

i need to review this stuff a bit :(

OpenStudy (anonymous):

is it ok if i ask you for help tomorrow?

OpenStudy (anonymous):

yeah of course

OpenStudy (anonymous):

No one likes perms and combs grrrrrrr

OpenStudy (anonymous):

yep...I hated this stuff when I did it years ago and never looked back

OpenStudy (anonymous):

but u seem really good with this anyways :)

OpenStudy (anonymous):

i ll go study other stufff now. ciao ciao :)